OpenStudy (anonymous):

Help? Tried to do something with the Compound Interest Formula but hit a wall? Rae owes $2,019 on a credit card with a 17.7% interest rate compounded monthly. What monthly payment should she make to pay off this debt in three years, assuming she does not charge any more purchases with the card?

OpenStudy (anonymous):

This is what i have so far. Compound interest formula: A = P(1 + r/n)^nt A = 2019(1 + 0.177/12)^12(3) A = 2019(1 + 0.177/12)^36 A = 2019(1 + 0.01475)^36 A = 2019(1.01475)^36 A = 2019(1.69404974724) A = 3420.29 I was thinking i could do that and divide the answer by 36 (cause thats how many months are in 3 years) but its not one of my answer choices...

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the formula you are using is for compound interest, not payments the formula for payments in entirely different

OpenStudy (anonymous):

Oh ok. What is the formula for payments? @misty1212

OpenStudy (misty1212):

\[\huge \frac{P\times \frac{apr}{n}}{1-(1+\frac{apr}{n})^{-ny}}\]

OpenStudy (misty1212):

in your case \(n=12,y=3, P=2019,r=.177\)

OpenStudy (misty1212):

\[\huge \frac{2019\times \frac{.177}{12}}{1-(1+\frac{.177}{12})^{-36}}\]

OpenStudy (anonymous):

Wow thanks! Ill let you know what i get. :)

OpenStudy (misty1212):

monthly payments about \($72.69\) http://www.wolframalpha.com/input/?i=+ \frac{2019\times+\frac{.177}{12}}{1-%281%2B\frac{.177}{12}%29^{-36}}

OpenStudy (anonymous):

Awesome! Thank you so much! :)

OpenStudy (anonymous):

I think i did something wrong... Can someone check my math??